OpenStudy (anonymous):

HELP PLEASE!!!! solve the inequality y + 6 ≥ 9

OpenStudy (acxbox22):

isolate y by subtracting 6 on both sides

sammixboo (sammixboo):

All you need to do is subtract 6 on both sides of the equation y + 6 - 6 >= 9 - 6 y >= 9 - 6 What's 9 - 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Answer is 3

OpenStudy (anonymous):

the answers are:

OpenStudy (acxbox22):

so now we have y>=3 <<this is our simplified inequality

sammixboo (sammixboo):

>= is the same as \(\ge\)

OpenStudy (anonymous):

thanks:)